water was added to 90g of a salt xcl2 to produce 40cm3 of a saturated solution at 25 o C . the solubility of the salt at 25 o C is 8 mol dm-3. if the atomic mass is 29 calculate the mass of the undissolved salt
1
Expert's answer
2020-06-05T05:13:00-0400

Solubility=8mol/10cm3=0.8mol/cm3

Yeild=40cm3

Then the mass that dissolved is

"Mass= moles * rfm"

Mass=0.8*29= 23.2g

The mass that did not dissolve becomes ; 90-23.2=66.8g